OpenStudy (anonymous):

Write the equation of a line with slope m = 3 and including point (3, 5).

OpenStudy (anonymous):

Just use the point slope form, where m=3, \((x_1,y_1)\) \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

Like this, \[y-5=3(x-3)\]

OpenStudy (anonymous):

Expand the RHS y-5=3x-9 Add 5 to both sides to isolate y, and you're done

OpenStudy (anonymous):

How do you "isolate" y?